# Benchmarking ECG Foundational Models: A Reality Check Across Clinical Tasks

### v1

@@ A.2.1 ECG-CPC model
The model architecture largely follows (Mehari & Strodthoff, 2023) and is composed of four convolutional layers as the encoder, followed by four S4 layers (Gu et al., 2022) (state dimension 8 and model dimension 512) as the predictor. In contrast to (Mehari & Strodthoff, 2023), the model operates at a sampling frequency of 240 Hz, which is the minimal sampling frequency in the HEEDB dataset (Koscova et al., 2024) used for pretraining. To account for the deviation from the sampling frequency in the original publication, the model uses a kernel size of 3 and a stride of 2 in the first convolutional layer and predicts ahead 14 steps (as compared to 12 originally) in the CPC objective.
@@ A.2.2 S4 model
The S4-based supervised baseline follows the specification in (Strodthoff et al., 2024). It also uses four S4 layers (Gu et al., 2022) (state dimension 8 and model dimension 512) without a convolutional encoder. It operates at a sampling frequency of 100 Hz with an input size of 2.5 s.
@@ A.3.1 Finetuning
Foundation Models (Finetuned) Supervised
@@ A.3.2 Frozen evaluation
Foundation Models (Frozen Evaluation) Supervised
@@ A.3.3 Linear evaluation
Foundation Models (Linear Evaluation) Supervised
@@ A.3.4 Ranking
Table 7: Statistical ranking of foundation models across evaluation modes and datasets. Rankings (Finetuned/Frozen/Linear) are assigned based on statistical equivalence groups determined by bootstrap testing, where models not performing significantly worse than the best model share the same rank. Lower ranks indicate better performance. signifies evaluation of a model trained on another dataset (listed above).
Foundation Models (Finetuned/Frozen/Linear) Supervised
@@
Table 8: Median statistical rankings of foundation models across evaluation modes by categories. Rankings (Finetuned/Frozen/Linear) represent the median performance position across all datasets within each category. Lower values indicate better overall performance.
Foundation Models (Finetuned/Frozen/Linear) Supervised
@@ A.4 Label-specific model predictions
Label ECGFounder ECGJEPA ECGFMKED MERL STMEM Hubert ECGFM ECGCPC S4 Net1D
@@
Label Units ECGFounder ECGJEPA ECGFMKED MERL STMEM Hubert ECGFM ECGCPC S4 Net1D Baseline
@@
Label Units ECGFounder ECGJEPA ECGFMKED MERL STMEM Hubert ECGFM ECGCPC S4 Net1D Baseline
@@
Label Units ECGFounder ECGJEPA ECGFMKED MERL STMEM Hubert ECGFM ECGCPC S4 Net1D Baseline
@@
Label Units ECGFounder ECGJEPA ECGFMKED MERL STMEM Hubert ECGFM ECGCPC S4 Net1D Baseline
@@
Label Units ECGFounder ECGJEPA ECGFMKED MERL STMEM Hubert ECGFM ECGCPC S4 Net1D Baseline

### v2

@@ A.2.1 ECG-CPC model
The model architecture largely follows (Mehari and Strodthoff, 2023) and is composed of four convolutional layers as the encoder, followed by four S4 layers (Gu et al., 2022) (state dimension 8 and model dimension 512) as the predictor. In contrast to (Mehari and Strodthoff, 2023), the model operates at a sampling frequency of 240 Hz, which is the minimal sampling frequency in the HEEDB dataset (Koscova et al., 2024) used for pretraining. To account for the deviation from the sampling frequency in the original publication, the model uses a kernel size of 3 and a stride of 2 in the first convolutional layer and predicts ahead 14 steps (as compared to 12 originally) in the CPC objective.
@@ A.2.2 S4 model
The S4-based supervised baseline follows the specification in (Strodthoff et al., 2024). It also uses four S4 layers (Gu et al., 2022) (state dimension 8 and model dimension 512) without a convolutional encoder. It operates at a sampling frequency of 100 Hz with an input size of 2.5 seconds.
@@ A.3.1 Finetuning
FMs (Finetuned) Supervised
@@ A.3.2 Frozen evaluation
FMs (Frozen Evaluation) Supervised
@@ A.3.3 Linear evaluation
FMs (Linear Evaluation) Supervised
@@ A.3.4 Ranking
Table 6: Statistical ranking of FMs across evaluation modes and datasets. Rankings (Finetuned/Frozen/Linear) are assigned based on statistical equivalence groups determined by bootstrap testing, where models not performing significantly worse than the best model share the same rank. Lower ranks indicate better performance. signifies evaluation of a model trained on another dataset (listed above).
FMs (Finetuned/Frozen/Linear) Supervised
@@
Table 7: Median statistical rankings of FMs across evaluation modes by categories. Rankings (Finetuned/Frozen/Linear) represent the median performance position across all datasets within each category. Lower values indicate better overall performance.
FMs (Finetuned/Frozen/Linear) Supervised
@@ A.4 Label-specific model predictions
Label Units ECGFounder ECG-JEPA ECGFM-KED MERL ST-MEM HuBERT-ECG ECG-FM ECG-CPC S4 Net1D Baseline
@@
Label Units ECGFounder ECG-JEPA ECGFM-KED MERL ST-MEM HuBERT-ECG ECG-FM ECG-CPC S4 Net1D Baseline
@@
Label Units ECGFounder ECG-JEPA ECGFM-KED MERL ST-MEM HuBERT-ECG ECG-FM ECG-CPC S4 Net1D Baseline
@@
Label Units ECGFounder ECG-JEPA ECGFM-KED MERL ST-MEM HuBERT-ECG ECG-FM ECG-CPC S4 Net1D Baseline
@@
Label Units ECGFounder ECG-JEPA ECGFM-KED MERL ST-MEM HuBERT-ECG ECG-FM ECG-CPC S4 Net1D Baseline
@@ A.5 Scaling analysis
A.6 Model complexity
In Table 32, we compare different notions of model complexity, namely, number of model parameters, number of floating point operations, peak GPU memory usage and inference throughput.
Table 32: Unified comparison of computational cost, memory usage, and inference efficiency for all ECG representation learning models. Reported metrics include: (i) GFLOPs for forward (F) and backward (B) passes (lower is better), measured with batch size 1 on an NVIDIA L40; (ii) peak GPU memory during inference (lower is better), measured on PTB-XL (all) with batch size 64; and (iii) throughput (samples/s; higher is better) and latency (ms/sample; lower is better) under the same hardware and batch size. Timesteps correspond to the input sequence length for each model, and parameter counts include all trainable weights. The three best, i.e., most efficient, FMs are marked in bold face. The overall best model is additionally underlined.
[…]
ECGFounder (CNN) 1250 33.8M 0.602 / 5.066 187.37 2220.67 / 0.450
ECG-JEPA (Transformer) 2500 87.2M 73.877 / 221.6 2136.79 98.71 / 10.131
ST-MEM (Transformer) 600 90.3M 20.926 / 62.779 609.71 368.64 / 2.713
MERL (CNN) 1250 4.6M 0.863 / 2.582 86.24 2727.36 / 0.367
ECGFM-KED (CNN) 5000 9.7M 5.423 / 16.258 427.39 1631.28 / 0.613
HuBERT-ECG (Transformer) 500 97.2M 18.829 / 69.779 845.27 213.87 / 4.676
ECG-FM (Transformer) 2500 93.9M 29.242 / 109.794 1091.71 692.35 / 1.444
ECG-CPC (SSM) 600 3.8M 1.741 / 5.213 482.32 442.38 / 2.260
A.7 Impact of layer-dependent learning rates during finetuning
In Table 33, we investigate the effect of layer-dependent learning rates during finetuning across three datasets, PTB-XL, EchoNext and CPSC-Extra.
Table 33: Impact of layer-dependent learning rates on model performance. AUROC scores comparing models trained with and without layer-dependent learning rate scheduling. Layer-dependent learning rates apply different learning rates across network layers, typically with lower rates for earlier layers. Higher AUROC values indicate better performance. Best scores are marked in bold face. Underlined values indicate models pretrained on the corresponding dataset.
Model PTB-XL (all) EchoNext CPSC-Extra
With $\uparrow$ / Without $\uparrow$ With $\uparrow$ / Without $\uparrow$ With $\uparrow$ / Without $\uparrow$
ECGFounder (CNN) 0.934 / 0.936 0.817 / 0.818 0.906 / 0.907
ECG-JEPA (Transformer) 0.940 / 0.779 0.817 / 0.730 0.897 / 0.790
ST-MEM (Transformer) 0.908 / 0.891 0.816 / 0.751 0.883 / 0.853
MERL (CNN) 0.925 / 0.919 0.822 / 0.823 0.873 / 0.852
ECGFM-KED (CNN) 0.889 / 0.918 0.806 / 0.817 0.824 / 0.833
HuBERT-ECG (Transformer) 0.915 / 0.499 0.792 / 0.500 0.876 / 0.513
ECG-FM (Transformer) 0.927 / 0.504 0.772 / 0.499 0.862 / 0.517
ECG-CPC (SSM) 0.949 / 0.938 0.831 / 0.828 0.898 / 0.875
A.8 Backbone comparison for supervised training from scratch
In Table 33, we provide additional results on the relative performance of different model backbone across three datasets, PTB-XL, EchoNext and CPSC-Extra. In addition to the S4-model, a structured state space model, and Net1D, a CNN, we consider Mamba-1 (Gu and Dao, 2024) and Mamba-2 (Dao and Gu, 2024) as an examples for more recent state-space model. More specifically, for the two Mamba models we use the following configuration d_model = 512, d_state = 128, n_layers = 4. The S4 backbone performs best or consistent with the best-performing model across all three datasets. Net1D is competitive only on a single dataset. Mamba-1 is the best perfoming model on a single dataset. This underscores the superiority of the S4-based baseline, in line with recent claims in the literature (Somvanshi et al., 2025; Patro and Agneeswaran, 2025) putting into question the suitability of latest state space models for the processing of physiological time series.
Table 34: Comparison of aggregated macro-AUROC scores for different model architectures trained from scratch. Higher AUROC values indicate better performance. The best-performing result is highlighted in boldface and underlined, while models that do not perform statistically significantly worse are also highlighted in boldface.
Model PTB-XL (all) $\uparrow$ EchoNext $\uparrow$ CPSC-Extra $\uparrow$
S4 (SSM) 0.941 0.819 0.852
Net1D (CNN) 0.929 0.818 0.803
Mamba-1 (SSM) 0.895 0.770 0.858
Mamba-2 (SSM) 0.914 0.807 0.844
A.9 Representational analysis based on CKA
A.9.1 Intra-model analysis
We start by describing the intra-model CKA patterns in detail. We investigate the FMs in their original form without any task-specific finetuning. The results are compiled in Figure 4.
ECG-CPC In Figure 4(a), the early CNN layers (CNN0-CNN2) show a very high similarity suggesting that they learn very similar representations, which could be taken as a sign that the CNN encoder is potentially overparameterized for the task. The final CNN layer (CNN3) shows similarity with both the earlier CNN layer as well as the first SSM layer (S4-0) thereby serving as a transition point between the local, CNN-based feature extractor and the sequential SSM layers in the hybrid architecture. The still pronounced dissimilarity between the the final CNN layer (CNN3) and the first SSM layer (S4-0) can most likely be explained by the difference in model architecture between these two layers. The following S4 layers become increasingly dissimilar to the CNN encoder layers, suggesting a substantial evolution throughout the architecture. Finally, the last S4 layer (S4-3) is highly specialized and shows only moderate to low similarity to other S4 layers, suggesting it learns the most task-specific representations for the CPC task. Overall the model shows a clear representational gradient from low-level CNN-features to high-level sequential features.
ECGFounder In Figure 4(b), we show the CKA analysis for the ECGFounder model. The convolutional stem shows low similarity with all other layers. Apart from that, the CNN model shows a gradual transition across layers with no abrupt transitions as observed in the hybrid ECG-CPC model between the convolutional feature extractor and the S4 layers. The model shows highly similar representations between S0 and S4, which suggests that these layers learn quite similar representation, suggesting a potential overparameterization in the middle of the network. The final two layers (S5, S6) show less similarity with earlier layer, suggesting that these are used for task specialization.
ECG-JEPA Figure 4(c) shows the CKA analysis for the ECG-JEPA transformer model. Again the embedding layer shows little similarity to all other layers. The first transformer layer (Blk0) serves a transition point between the embedding layer and the following transformer layers, showing moderate similarity to both. As a striking feature, the following transformer layers (Blk1-Blk10) shows a very high similarity, which suggests that the representational structure barely changes across the 10 layers. This might be indicative for a representation collapse and possible inefficiency in the architecture where most layer contribute only minimally distinct representations. The representations of the final transformer layer (Blk11) deviate strongly from the previous, suggesting that the former is used for task specialization.
ST-MEM In Figure 4(d), we see a similar overall structure as for the also transformer-based ECG-JEPA model. The embedding layer and the first transformer layer are distinct from a largely homogeneous block of following layers (Blk1-Blk9), the final layers (Blk10-Blk11) specialize. Compared to ECG-JEPA, ST-MEM shows a slightly better gradient through the middle layers, still the middle part of the architecture remains quite homogeneous, which again suggests wasted capacity. A remarkable difference compared to ECG-JEPA is the low similarity of the first transformer block (Blk0) with the following blocks (Blk1 onwards). This suggests a representational bottleneck early in the network which prevents the propagation of information from the embeddings into the following transformer blocks.
A.9.2 Inter-model analysis
(a) Early stage
(b) Mid stage
(c) Late stage
Figure 5: Inter-model representational similarity across network depths. CKA heatmaps comparing corresponding stages across four FMs (ECGFounder, ECG-JEPA, ST-MEM, ECG-CPC). Higher values (yellow) indicate similar representations between layers. CKA computed using Gaussian RBF kernel ($\sigma=1.0$) on 2,500 samples of PTB-XL (all) dataset per model.
In Figure 5, we analyze the representational similarity across different models at three different points in the model architectures (early: after the encoder/stem/tokenizer; mid: after the first half of the processing layers; late: after the final processing layer).
In the early representations, shown in Figure 5(a), the two transformers show high mutual similarity while the two non-transformer models show other patterns. This might be taken as a hint it is the architecture which shapes early representations and not the training objective. ECGFounder’s early features are most dissimilar to all other features, indicating that the pure CNN learns qualitatively different early features.
This result should be compared to the CKA comparison of the late features, which is shown in Figure 5(c). The two transformer models (ECG-JEPA and ST-MEM) remain highly similar, whereas ECGFounder remains very dissimilar to all three other models. Also the dissimilarity of ECG-CPC and the transformer models increases in later layers, which supports the hypothesis that the ECG-CPC model learns distinct representations and does not replicate the patterns learned by the transformer models.
A.9.3 Insights
To summarize, in the hybrid CNN-SSM model (ECG-CPC) the CNN encoder extracts local features while the S4 layers progressively build temporal abstractions. Each component serves a distinct purpose, and you see representational change at each layer.
The CNN-based ECGFounder shows smooth gradients, but still high similarities across blocks S0-S4 indicate redundancies, which could possibly alleviated by decreasing the number of layers. Both considered transformer models show nearly identical representations in the middle 8-10 layers, which puts into question the efficiency of the archtitecture. It raises the question why transformer models show this particular pattern on ECG data specifically. The isolated Blk0 in the case of ST-MEM indicate poor integration of encoder and further processing layers, which might explain the deficiencies in downstream performance observed for this model.
The inter model comparisons reveal strong similarities between the two transformer models in spite of different training objectives. The results indicate potential for ensembling of ECGFounder with any of the other models to maximize diversity.
The CKA analysis provided hints for ST-MEM’s weakness (early bottleneck) and ECG-CPC’s strengths (efficient, purposeful layer utilization), which establishes CKA as a useful diagnostic tool for ECG FMs.
A.10 AUROC vs. Specificity vs. F1-Score
In this section, we shed light on alternative performance metrics beyond macro AUROC. The clinically most commonly used performance metrics such as sensitivity and specificity require the specification of a decision threshold. While this decision should be taken based on clinical desiderata in a disease-specific manner, we restrict for simplicity to a disease-unspecific threshold. More specifically, we adjust the decision threshold for every condition such a sensitivity of 0.8 is reached. We then assess model performance according to macro-averaged specificity and F1-score. Due to its widespread use in the literature, we base the experiment on the PTB-XL on the most finegrained level of the label hierarchy covering 71 unique labels. We compare the performance of foundation models after finetuning. The results are compiled in Figure 6.
As a non-trivial finding, the model ranking according to AUROC completely agrees with the model ranking according to specificity, as it suggests that better overall discriminative power according to AUROC translates into better model performance in terms of clinically relevant performance metrics sensitivity/specificity at relevant operating points. The F1-score assesses only precision/positive predictive value as the recall/sensitivity was fixed to 0.8 for all models. The model ranking based on F1-score shows correlation with the ranking based on AUROC but does not fully agree. We see the agreement in terms of specificity as the more crucial observation as the positive predictive value depends on the prevalence of the respective conditions in the dataset, which, however, does not follow a proper population-level distribution.
(a) AUROC
(b) Specificity
(c) F1-score
Figure 6: Performance comparison of foundation models on PTB-XL under the finetuning with linear head evaluation mode. Evaluated using (a) AUROC, (b) Specificity, and (c) F1-score at target sensitivity of 0.8. Higher values indicate better performance.